Calculate the density of a mixture of gases that has an average molar mass of 26.7 g/mol at a temperature of 32.1 C and 678 mmHg.
1
Expert's answer
2020-11-12T07:10:53-0500

PV = nRT P = 678mmHg = 90392.6Pa, T = 32.1°C = 305.1K

We assume the number of mole is 1

V = nRT/P

V = 1×8.314×305.1/ 909392.6

V= 0.028m3


Density = Mass/volume

= 26.7 × 10-3 Kg / 0.028m3

= 0.95Kgm-3